# MapGPT: Map-Guided Prompting for Unified Vision-and-Language Navigation

### 14 Jan 2024 v1

@@ 4.3 Ablation Study
As shown in Table 4, we conduct an ablation study to explore the effectiveness of some crucial information in the prompt on 72 scenes of the R2R dataset. For basic prompt systems that do not employ map and planning, we discover that the extraction of visual information significantly impacts navigation experts. When we only include scene-level description, i.e., a caption, for the visual observation, our baseline can only obtain a success rate of 27.8%. However, with the inclusion of object detection, there is a significant improvement in success rate, reaching 32.9%, while the OSR remains unchanged. This could be attributed to the fact that the results of object detection assist the agent in better identifying the correct locations to stop, while providing limited assistance in the process of exploration.
Based on this baseline system, we first introduce map information, which significantly enhances the agent’s global exploration capability and improves the accuracy in OSR by over 10%. These performance gains also translated partially into the final task success rate, increasing it from 32.9% to 37.5%. Furthermore, we incorporate a planning iterative updating mechanism, enabling the agent to reference path planning from the last step. This mechanism does not significantly improve the OSR performance since the map already provides sufficient capability for global exploration. However, it influences the process of decision-making, raising the success rate from 37.5% to 41.2% ultimately.
@@ 4.4 Case Study
In Figure 3, we showcase an example from REVERIE that sufficiently demonstrates the various abilities brought to the agent benefiting from map-guided prompting and iterative updating of multi-step path planning. In step 2, the agent is located at Place 1. Benefiting from the map understanding, it astutely identifies that out of the five places connected to its current position, only three remain unexplored. Since all these three places are described as bedrooms, the agent makes a multi-step path planning to sequentially explore them to find the bathroom. By Step 4, the agent has successfully executed the previous planning and all planned places except Place 10 have been explored. However, as Place 10 is not accessible from Place 9, the agent demonstrates its global exploration based on the map and decides to backtrack to Place 1 to continue exploring Place 10. The path planning information is also transmitted to subsequent navigation steps through an iterative updating mechanism. The latest planning is “go back to Place 1 and then explore Place 10…”, while the previous planning at step 2 is “I will first go to Place 8, then Place 9, and finally Place 10…”. This avoids the interference caused by recording all thoughts of each step and ensuring accurate execution of multi-step path planning. If the agent encounters new situations, it can also autonomously decide to remake a new path planning.
@@ 5 Conclusion
In this paper, we propose a novel zero-shot agent, named MapGPT, for the VLN task. MapGPT utilizes map-guided prompts to activate the global exploration and path-planning capabilities of the LLMs, thus avoiding getting stuck in the local environment. Additionally, we require the agent to explicitly generate and iteratively update planning to further enhance navigation performance. Extensive experiments demonstrate the impressive performance of MapGPT on both the R2R and REVERIE datasets.
@@ A.1 Prompts
We provide specific task description prompts that are directly fed into the system content of GPT [27, 8] API. As shown in Figure 4, our unified prompts consist of three parts, namely task background, input definitions, and output requirements.
@@
We have achieved effortless adaptation between fine-grained R2R [5] instructions and high-level REVERIE [32] instructions with only a few intuitive and necessary modifications. These modifications are primarily utilized for ignoring the extensive interactive actions with objects in REVERIE, since our unified agent is designed to focus on the navigation task. The stopping conditions also differ, as R2R only requires the agent to stop at the destination, while REVERIE demands the agent to check the target object in its surroundings before stopping. Additionally, unlike R2R, REVERIE instructions contain a significant amount of details regarding landmarks and objects. Due to the limitations of the visual model’s capability, we are unable to fully acquire this information. Therefore, we modify the prompts to require the agent to ignore these details.
A.2 Implementation
We implement the MapGPT agent using the GPT [27, 28, 29] API and the “gpt-4-0613” version of the model is adopted for GPT-4-based MapGPT. The “gpt-3.5-turbo-0613” and “gpt-4-vision-preview” version of the models are employed for GPT-3.5-based and GPT-4V-based MapGPT, respectively. Additionally, we set the temperature parameter of GPT models to 0, and all experiments are performed with a single-prediction setting. Moreover, we employ the visual model BLIP-2 [23] to extract captions as scene descriptions, using the released “blip2-flan-t5-xl” model. It is worth noting that we utilize identical prompts for GPT-3.5 and GPT-4 models. As for GPT-4V, we make no additional modifications, only incorporate some descriptions of image modality and enable the agent to directly reference the provided images corresponding to places in the action space and history, eliminating the need for converting images into textual descriptions.
Appendix B Additional Experimental Results
[…]
MapGPT 5.80 61.6 41.2 25.4
   w/o Trajectory 6.34 59.7 38.4 23.8
   w/o Supplementary Info 6.52 59.3 37.5 24.2
   w/o Updating Mechanism 6.43 58.3 35.6 21.8
Table 5: Ablation on the R2R dataset about removing “Trajectory” or “Supplementary Info” in the map prompts, as well as the iterative updating mechanism of the multi-step planning.
In the main text, we have reported the results of removing all the map-related or planning-related prompts. Here, we conduct further ablation study on some design details of prompts related to maps and planning.
As shown in Table 5, We retain the core prompts of the map and planning, while separately removing the “Trajectory” or “Supplementary Info” from the map, as well as the iterative updating mechanism of multi-step planning. The trajectories assist the agent in understanding its navigation path within the map, while the supplemental information represents descriptions of nodes that are currently inaccessible. Removing these pieces of information affects the agent’s perception and decision-making regarding the environment, subsequently impacting its performance. Therefore, these prompt designs are necessary.
We also test requiring the agent to output only some multi-step planning at each step, which can be considered as another form of “thought”, instead of using an iterative updating mechanism to propagate planning information to the next step. It can be observed that this prompt design is harmful to navigation performance, as the agent engages in redundant long-term planning that is not utilized in the following steps. Therefore, it is necessary to iteratively update and guide the agent’s subsequent decisions based on multi-step planning.
Appendix C More Qualitative Examples
We provide additional successful and failure cases on the R2R and REVERIE datasets to qualitatively analyze the capabilities and limitations of our proposed MapGPT (GPT-4 based).
Figure 5 demonstrates a successful case on the R2R dataset. In step 6, after thoroughly exploring places 3 and 4 connected to place 1, the agent decides to backtrack to place 1 and subsequently explore currently inaccessible places 6 and 7. Ultimately, the agent successfully terminates at place 7 in step 10. Figure 6 shows another successful case on the REVERIE dataset. This example shares similar scenarios with the example in the main text, but it differs in terms of the starting point and the given instruction. Here, we present a more detailed navigation process. In step 4, the agent makes appropriate planning based on the map and sequentially explores places 11, 12, and 13 in search of a bathroom. In step 8, after exploring place 13, the agent discovers that place 14, which is connected to the current location, contains a toilet and sink, indicating a potential correct direction. Therefore, the agent continues exploring and finally stops at the correct destination.
As shown in Figure 7, we further summarize three typical types of failure cases. (a) The process of converting visual observations into textual descriptions may result in error or information loss. For example, the agent recognizes the room as a hallway, causing the agent to fail to stop at the correct location. (b) The agent may fail to follow the details in the instructions accurately. For instance, instead of walking straight into a bedroom in the eleven o’clock direction as instructed, it turns left in step 1 and enters another incorrect bedroom, and stops there. (c) The scenes are highly challenging, and the instructions may not provide many clues. Thus, the agent may fail to explore the correct direction in time. To address these limitations, one possible research direction for the future is to utilize increasingly powerful multimodal models and combine them with map-guided prompts for making multi-step planning directly, without the need to convert visual observations into textual descriptions.
@@ Appendix C More Qualitative Examples
Figure 4: Task description prompts for the R2R and REVERIE datasets. We make some simple yet necessary modifications to transfer the MapGPT from the R2R task to REVERIE. This work focuses on unified navigation, while instructions in REVERIE often require some interactive actions on objects. Therefore, we require the agent to ignore these actions, as well as various object details that are difficult for the vision models to extract adequately.
@@
Figure 6: A successful example on the REVERIE dataset. We demonstrate some crucial steps that leverage map-guided global exploration and planning capabilities, ultimately resulting in successful navigation.
Figure 7: Three typical reasons for failure. (a) The error or loss of information due to the transformation from visual observation to textual description (the agent passes the correct location but fails to stop). (b) Stopping erroneously in similar locations. (c) Failure to timely backtrack and explore the correct direction if the scene is complex and the instruction does not provide much guidance (the agent has explored place 1 and plans to explore place 2, while the correct direction is 0$\rightarrow$3).

### 25 Feb 2024 v2

@@ 4.3 Ablation Study
As shown in Table 4, we explore the effectiveness of different prompt designs on 72 various scenes of the R2R dataset.
Coordinate Maps vs. Topological Maps. DUET encodes the precise GPS coordinates of each node for graph learning. We also attempt to input these coordinates into GPT for understanding and reasoning. However, the agent appears to struggle with global exploration based on the coordinates, resulting in a performance decline compared to the baseline that does not utilize the map. On the other hand, our proposed map-guided prompting method provides the topological relationship in a natural language form for GPT-4 and GPT-4V to understand, leading to a significant performance improvement.
Global Action Planning vs. Adaptive Path Planning. DUET develops a global action planning, which involves selecting an optimal node from both accessible and unexplored inaccessible nodes, and teleporting the agent to that node. We also implement a similar agent that incorporates nodes from the “Supplementary Info" into the action space for action planning by GPT-4V. Experimental results indicate that this approach does not significantly improve the zero-shot agent equipped with GPT. Instead, we adopt an adaptive planning approach, where GPT explicitly outputs a segment of multi-step path planning, allowing flexible attention to multiple potential nodes or sub-goals and the ability to correct previous errors. This effectively leverages the advantages of the GPT for thinking and planning, rather than selecting a single action.
@@ 4.4 Case Study
In Figure 3, we showcase a successful example from REVERIE that demonstrates the various abilities of MapGPT (GPT-4V based). This example poses some challenges for a zero-shot agent since the places observed at place 4, namely places 5 and 9, as well as the previously observed places 1, 6, 7, and 8, do not contain the bathroom which is behind the door. Additionally, the instructions in REVERIE typically do not include information about turning or any other specific actions. Therefore, the agent needs to explore the entire environment to determine the correct direction. Benefiting from map-guided prompting with adaptive planning, MapGPT demonstrates a strong understanding of the topological relationships between nodes and adaptively performs multi-step path planning. Based on six unexplored global candidates, the agent systematically conducts global exploration by selecting the four most probable nodes, as they are situated within the bedroom and are more likely to be connected to the bathroom. Besides, the planning content mentions the possibility of backtracking to place 2 for re-exploration if necessary, and is also adaptively updated upon discovering the direct connection between places 5 and 4 in the map. After several steps, when the agent moves to place 8, it discovers the bathroom hidden behind the door and successfully reaches the destination.
@@ 5 Conclusion
In this paper, we propose a novel zero-shot agent, named MapGPT, for the VLN task. MapGPT utilizes map-guided prompting, which builds online constructed maps using prompts that provide GPT with node information and topological relationships to activate global exploration. Additionally, we propose an adaptive planning mechanism that enables multi-step path planning based on the map, allowing the agent to systematically explore potential objectives. Through extensive experiments, we demonstrate that MapGPT achieves state-of-the-art zero-shot performance with global thinking and path planning capabilities.
Limitations
Despite the significant performance gap between MapGPT and models based on pre-training and fine-tuning, zero-shot VLN still holds significant research value. It requires no data training, thus demonstrating great potential in terms of generalization and sim-to-real transfer, as GPT’s pre-training corpus contains a large amount of real-world image data, while data from simulators is comparatively limited. However, MapGPT is only experimented within a simulator that incorporates certain ideal assumptions, which is a limitation. Developing GPT-based agents directly in the real world and addressing various real-world challenges would be a meaningful future direction.
@@ A.1 Prompts
We provide specific task description prompts that are directly fed into the system content of GPT-4V OpenAI (2023b, c) API. As shown in Figure 4, our unified prompts consist of three parts, namely task background, input definitions, and output requirements.
@@
We have achieved effortless adaptation between fine-grained R2R Anderson et al. (2018b) instructions and high-level REVERIE Qi et al. (2020b) instructions with only a few intuitive and necessary modifications. These modifications are primarily utilized for ignoring the extensive interactive actions with objects in REVERIE, since our unified agent is designed to focus on the navigation task. The stopping conditions also differ, as R2R only requires the agent to stop at the destination, while REVERIE demands the agent to check the target object in its surroundings before stopping.
Templates
For a two-stage prompt system, we formulate the observations that have been converted into text as $O_{t}=\{o_{t,i}\}_{i=1}^{N}$, where $N$ is the number of navigable viewpoints at step $t$. Each $o_{t,i}$ represents the observation towards a specific navigable point and is formulated using the prompt template:
“<{scene}>, which also includes <{objects}>”.
Besides, we have defined six directional concepts, namely “go forward to”, “turn left to”, “turn right to”, “turn around to”, “go up to”, and “go down to”, according to the directions of navigable viewpoints. Thus, a complete action $a_{t}$ could be “B. turn around to Place 1: <a room with blue walls>, which also includes <bed, curtain, picture>”.
For a one-stage system based on GPT-4V, we directly require the agent to refer to images corresponding to various places. At the end of all prompts, we input the image IDs and original images in an interleaved format, such as “Image 0: <Img0> Image 1: <Img1> …”. When used in the action space, an action $a_{t}$ could be “C. turn around to Place 2 which is corresponding to Image 2”. Prompts in history and maps also employ similar templates.
A.2 Implementation Details
We have built the first GPT-4V-based OpenAI (2023b, c) agent in the VLN field, directly processing multimodal inputs in one stage. For a fair comparison with previous methods specifically designed for GPT-4, we have also implemented a two-stage system, where we follow NavGPT Zhou et al. (2023) and utilize BLIP-2 Li et al. (2023) to provide a caption for the observation, and employ Faster R-CNN Ren et al. (2015) to detect existing objects. Our core contributions, namely map-guided prompting and adaptive path planning, can be applied to both of these systems. To adapt our MapGPT to the REVERIE dataset, we only make some simple yet necessary modifications, which demonstrates that our MapGPT is more unified in the VLN task.
@@ A.2 Implementation Details
Figure 4: Task description prompts for the R2R and REVERIE datasets. We make some simple yet necessary modifications to transfer MapGPT from the R2R task to REVERIE. This work focuses on unified navigation, while instructions in REVERIE often require some interactive actions on objects. Therefore, we require the agent to ignore these actions.
Appendix B More Qualitative Examples
We provide additional successful and failure cases to qualitatively analyze the capabilities and limitations of our proposed MapGPT.
Figure 5 demonstrates a successful case on the R2R dataset. In step 6, after thoroughly exploring places 3 and 4 connected to place 1, the agent decides to backtrack to place 1 and subsequently explore currently inaccessible places 6 and 7. Ultimately, the agent successfully terminates at place 7 in step 10.
As shown in Figure 6, we further summarize two typical types of failure cases, which are also common challenges for other zero-shot VLN agents. (a) The agent may fail to follow the details in the instructions accurately. For instance, instead of walking straight into a bedroom in the eleven o’clock direction as instructed, it turns left in step 1 and enters another incorrect bedroom, and stops there. (b) The scenes are highly challenging, and the instructions may not provide sufficient clues. Thus, the agent may fail to explore the correct direction in time.
@@ Appendix B More Qualitative Examples
Figure 6: Two typical reasons for failure. (a) Stopping erroneously in similar locations. (b) Failure to timely backtrack and explore the correct direction if the scene is complex and the instruction does not provide much guidance (the agent has explored place 1 and plans to explore place 2, while the correct direction is 0$\rightarrow$3).